# The Catalogue of Virtual Early-Type Galaxies from IllustrisTNG: Validation and Real Observation Consistency

### v1

@@ Abstract
Early-type galaxies (ETGs) are reference systems to understand galaxy formation and evolution processes. The physics of their collapse and internal dynamics are codified in well-known scaling relations. Cosmological hydrodynamical simulations play an important role, providing insights into the 3D distribution of matter and galaxy formation mechanisms, as well as validating methods to infer the properties of real objects. In this work, we present the closest-to-reality sample of ETGs from the IllustrisTNG100-1 simulation, dubbed "virtual-ETGs," based on an observational-like algorithm that combines standard projected and three-dimensional galaxy structural parameters. We extract 2D photometric information by projecting the galaxies’ light into three planes and modeling them via Sérsic profiles. Aperture velocity dispersions, corrected for softened central dynamics, are calculated along the line-of-sight orthogonal to the photometric projection plane. Central mass density profiles assume a power-law model, while 3D masses remain unmodified from the IllustrisTNG catalogue. The final catalogue includes $10121$ galaxies at redshifts $z\leq 0.1$. By comparing the virtual properties with observations, we find that the virtual-ETG scaling relations (e.g., size-mass, size-central surface brightness, and Faber-Jackson), central density slopes, and scaling relations among total density slopes and galaxy structural parameters are generally consistent with observations. We make the virtual-ETG publicly available for galaxy formation studies and plan to use this sample as a training set for machine learning tools to infer galaxy properties in future imaging and spectroscopic surveys.
@@ 1 Introduction
If, on one hand, photometry and spectroscopy serve as the cornerstone to the development of methods and models able to unveil the physics and structure of galaxies, on the other hand, they provide a limited view of these systems, as we can only observe a galaxy through its projection on the sky, leaving their three-dimensional shape and matter distribution accessible only via assumptions on their equilibrium and internal geometry. Here it is where hydrodynamical simulations (e.g., Schaye et al. 2014; Vogelsberger et al. 2014; Crain et al. 2015; Nelson et al. 2019; Wetzel et al. 2023) can play a crucial role, offering a powerful framework to explore galaxy formation in a cosmological context and validating the methods typically used to infer the internal structure of galaxies, including their dark matter content. e.g., Illustris simulations (Vogelsberger et al., 2014), have been used to validate dynamical models (e.g., the Jeans Anisotropic Modelling, Li et al. 2015).
Unfortunately, the process of solving the Jeans equations can be complex, often needing rather long computational times. A consequence is that the Jeans approach is more feasible for samples of thousands of galaxies (see, e.g., Tortora et al. 2014; Tortora & Napolitano 2022; Zhu et al. 2023b) while the upcoming large galaxy surveys will observe millions to billions of galaxies. In this context, machine Learning (ML) models present a promising avenue for addressing this challenge. One example of its use is as an alternative approach to inference of the total and dark matter content in galaxies (e.g., Von Marttens et al. 2022; Wu et al. 2024; Chu et al. 2024) or the emulation of dynamical models such as JAM with neural networks (Gomer et al., 2023). In particular, in Wu et al. (2024), we have provided a first application of these techniques to real samples and shown that machine learning can be as accurate as JAM or similar Jeans analyses, but providing total and dark masses of thousands of galaxies in a timescale of a few seconds.
@@ 2.1 The simulation
The galaxies are identified as gravitationally bound structures of stellar, gas and dark matter particles using the SUBFIND algorithm (Springel et al., 2001; Dolag et al., 2009). For this study, we selected only central galaxies that are the most massive objects identified by SUBFIND in each friend-of-friend (FoF) group and are of cosmological origin (SubhaloFlag=True), in the sense that they have formed due to the process of structure formation and collapse (Nelson et al., 2019). We filtered galaxies in nine snapshots corresponding to the redshift values $z=0,0.01,0.02,0.03,0.05,0.06,0.07,0.08,0.1$. We assume that in this redshift window there is no significant galactic evolution to impact the studies of this paper and, for now, we do not try to inspect the impact of redshift in the analysis. We also consider only the galaxies with total stellar masses in the range $10^{10.3}\leq M_{\star}/M_{\odot}\leq 10^{11.9}$, to guarantee a sufficiently large number of stellar particles per galaxy to minimize the Poissonian noise on the average quantities derived by the binning of the particles (see below). This first selection results in a sample of $23161$ objects for $z\leq 0.1$. Further filters to this sample (such as galaxy morphological classification) are discussed in detail in the next sections.
@@ 2.3.1 Stellar light measurement
where $\tau$ is the optical depth. In this model, the $\tau$ for young stellar systems is associated with dust in molecular clouds, in which the young stars are embedded. Following the disruption of molecular clouds, which occurs on a timescale of $\sim 10^{7}$ yr (Blitz & Shu, 1980), the optical depth is associated with a uniform screen across the galaxy. Following Charlot & Fall (2000), we use $\tau_{1}\sim 1.0$ and $\tau_{2}\sim 0.3$, that have been found to adequately describe a wide range of observations.
@@ 2.3.2 Galaxy geometry: ellipticity, centroid of luminosity, and orientation angle
Before we move to detail this procedure, it is important to recall that when a smaller halo is accreted onto a bigger structure, its dark and baryonic matter at the outskirts can be tidally stripped while sinking to the centre of the host. This stellar mass is still “weakly” gravitationally bound to the galaxy, and hence normally included in the raw measurement of SUBFIND. In the context of galaxy clusters, this stellar material is usually referred as intracluster light, and it generally accounts for a significant fraction of the total luminosity of a galaxy it overlaps with (Lin & Mohr, 2004; Zibetti et al., 2005; Puchwein et al., 2010). Despite its importance in the context of the galaxy and group/cluster formation theories, this component represents a contaminant of the individual galaxies, as it does not constitute a fully bound component of any specific individual galaxy, but rather a component eventually relaxing in the cluster potential (see, e.g., Pota et al. 2018; Spiniello et al. 2018), with a phase space distribution of the stars which is different from the individual galaxy populations. In simulations, in order to mitigate the impact of this component on the stellar mass and derived light profiles, a common technique is to make a 3D radial cut of $30$ kpc from the galactic centre (e.g., Pillepich et al. 2017; Lu et al. 2020), which is defined as the position of the particle with the minimum gravitational potential by the SUBFIND algorithm. In addition to this 3D cut, following Xu et al. (2017), for a given galaxy projection we define a 2D radial cut of $30$ kpc from the centre of light $(X_{gc},Y_{gc})$, defined as the average projected $(X_{i},Y_{i})$ position of each stellar particle weighted by their dust-attenuated luminosity value $L_{i}$:
@@
where $\sum_{i}L_{i}$ is the total luminosity within this $30$ kpc aperture. Given the values of $(X_{gc},Y_{gc})$, the measurement of the axis ratio $K=b/a$ and orientation angle $\phi$ of a galaxy were finally obtained through the computation of the second moments of luminosity, given by the set of equations:
@@ 2.3.3 Surface brightness profiles
As for the maximum radius to adopt to include stellar particles in the fit, we have used a standard 30 kpc sphere (see, e.g., Pillepich et al. 2018). We have also tried a more observationally motivated SB cut, corresponding to 26.5 mag/arcsec2 in $g$-band (see Tang et al. 2021, and also §1). However, since we have found no significant differences in the best-fit quantities that we will discuss later in this paper (in particular the effective radius and the total luminosity of the virtual-ETG sample), we have decided to discard the quantities based on the SB cut. The reason for the low sensitivity to the outer cut resides on the fact that this does not directly impact the inferred quantities that are always the result of a fitting procedure. This is different from the reason motivating the 26.5 mag/arcsec2 in $g$-band in Tang et al. (2021), who use as total quantities the ones obtained by integrating all particles within that limit, with neither fit or extrapolation.
@@
We chose the best-fit parameters based on a minimum $\chi^{2}$ approach, having used $R_{\rm e,dir}$ and $I_{\rm e,dir}$ for each galaxy as initial guesses for the fitting procedure, while $n=2.5$ is assumed as initial values for all galaxies. Using a least-squares algorithm [^6] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.least_squares.html, we fitted several profiles considering the different allowed values of $x$ (with $5$ steps between each sub-range) and chose the profile in which the fitted parameters give the minimum value of $\chi^{2}$. In short, ten profiles with different lower radial limits were fitted to each galaxy, and the best-fit was chosen based on the minimum $\chi^{2}$ value among them. After the fitting step, we took $R_{\rm e,mod}$, $I_{\rm e,mod}$ and $n$ as being the final half-light radius, half-light surface brightness and Sérsic index for each galaxy. For these ones, from now on, we drop the subscript mod, and refer to them as $R_{\rm e}$, $I_{\rm e}$ and $n$, respectively. Particular values of $n$ in Eq. (9) include $n=1$ that is the so-called exponential profile (Mo et al., 2010; Binney & Tremaine, 2011), and characterizes the surface brightness distribution of disky or other light-diffuse systems. For $n=4$ we have the De Vaucouleurs profile (De Vaucouleurs, 1948; Mo et al., 2010; Binney & Tremaine, 2011), that is better suited for the description of spheroidal systems, such as bulges and elliptical galaxies. In Fig. 1, we show the visual representations of two example galaxies and their respective projections. The one in the upper row has a low Sérsic index value ($n\lesssim 1$), and a disky appearance, which characterizes a Late-Type Galaxy (LTG). The galaxy in the lower row has a higher Sérsic index ($n\geq 4$) and has a smooth, featureless appearance, which are characteristics of an ETG. Furthermore, in Fig. 2 we present the surface brightness profiles for a pair of ETGs and a pair of LTGs, projected along the simulation’s $X$-axis. More details on the galaxy morphological classification are presented in §2.3.4.
@@ 2.6.1 Three-dimensional quantities
    Total 3D stellar mass, ($M_{\rm 3D,\star}^{\rm tot}$): this is simply derived as $M_{\rm 3D,\star}^{\rm tot}=M_{\rm tot}-M_{\rm DM}-M_{\rm gas}$. All these quantities are derived from the TNG100 SUBFIND catalogue777 For a projected total stellar mass definition, a straightforward approach is given by Eq. (12), that gives a logical way of estimating the modelled luminosity of a galaxy at any radius (Trujillo et al., 2001). Formally, the total luminosity of a galaxy is defined at an infinite extent, $r\to\infty$ in Eq. (12). Based on this total luminosity definition, the total stellar mass of a galaxy can be approximated as $M_{\star}(r\to\infty)=\Upsilon_{\star}L(r\to\infty)$. One should not expect this stellar mass definition to equalize the total stellar mass $M_{\rm 3D,\star}(<30~{}\text{kpc})$, as the latter is truncated at a finite radius while the former allows for an analytical extrapolation up to infinity. In fact, based on this extrapolation, $M_{\rm 2D,\star}^{\rm tot}(<\infty)$ should be greater than $M_{\rm 3D,\star}(<30~{}\text{kpc})$, specially at the bright end, where we would expect a greater impact of the intracluster light. With that in mind, we only discuss the comparison between the aperture luminosity and stellar mass definitions and their projected counterparts in §3.4.2 and §3.4.3..
@@ 2.6.2 Projected quantities
    where $\bar{v}$ is the mean velocity of the same $v_{i}$ sample defined above  [^9] We have also computed a light-weighted velocity dispersion, using the ratio of the particle luminosity over the total luminosity as weight, obtaining the same distribution of $\sigma_{[\epsilon_{\star},R_{e}]}$, consistently with the recent findings from Sohn et al. (2024), hence we decided to keep the simple Eq. (19).. This is the most unbiased LOS velocity dispersion we can “measure” in simulations that fully traces the internal dynamics of the simulated galaxies and that is closely connected to the “virial” velocity dispersion that one should include in the galaxy fundamental plane (de Graaff et al. 2022), although in Eq. (19) we include both random and rotational motions making this definition intrinsically larger than the rotation subtracted velocity dispersions derived from galaxy spectroscopy, especially toward the low mass end ($\log M_{\star}/M_{\odot}<11$) of our sample that is usually dominated by fast-rotators (see, e.g., Guo et al. 2020). Furthermore, real spectroscopy, yet differently from the $\sigma_{[\epsilon_{\star},R_{\rm e}]}$ above, considers all stars down to the very galaxy centres, where the velocity dispersion profile of ETGs, usually show their maximum value101010 For observations, the main “dilution” factor is the atmospheric seeing, which is generally ignored on fiber/long-slit observations, but it is accounted for in integral field spectroscopy (Chung et al. 2021 and references therein).. To take this central component, missed by Eq. (19), into account, we introduce a “corrected” $\sigma_{\rm e}$, which is fully detailed in Appendix D but that we also report here below for sake of completeness:
@@
    (see Appendix D for details). This $\sigma_{\rm e,corr}$ is the one we will use in the analysis and comparisons with literature, and will be included in the catalogue. Thus, from now on we will drop the subscript $\rm corr$ and refer to it just as $\sigma_{\rm e}$, unless it is needed to be specified.

### v2

@@ Abstract
Early-type galaxies (ETGs) are reference systems to understand galaxy formation and evolution processes. The physics of their formation and internal dynamics are codified in well-known scaling relations. In this context, cosmological hydrodynamical simulations play an important role in probing the physical origins of scaling relations by providing a controlled environment to study the formation and evolution of galaxies, linking their internal dynamics to underlying physical processes, and testing the robustness of observational inference methods. In this work, we present the closest-to-reality sample of ETGs from the IllustrisTNG100-1 simulation, dubbed "virtual-ETGs", based on an observational-like algorithm that combines standard projected and three-dimensional galaxy structural parameters. We extract 2D photometric information by projecting the galaxies’ light into three planes and modelling them via Sérsic profiles. Aperture velocity dispersions, corrected for softened central dynamics, are calculated along the line-of-sight orthogonal to the photometric projection plane. Central mass density profiles assume a power-law model, while 3D masses remain unmodified from the IllustrisTNG catalogue. The final catalogue includes $10121$ galaxies at redshifts $z\leq 0.1$. By comparing the virtual properties with observations, we find that the virtual-ETG scaling relations (e.g., size-mass, size-central surface brightness, and Faber-Jackson), central density slopes, and scaling relations among total density slopes and galaxy structural parameters are generally consistent with observations. We make the virtual-ETG publicly available for galaxy formation studies and plan to use this sample as a training set for machine learning tools to infer galaxy properties in future imaging and spectroscopic surveys.
@@ 1 Introduction
If, on one hand, photometry and spectroscopy serve as the cornerstone to the development of methods and models able to unveil the physics and structure of galaxies, on the other hand, they provide a limited view of these systems, as we can only observe a galaxy through its projection on the sky, leaving their three-dimensional shape and matter distribution accessible only via assumptions on their equilibrium and internal geometry. Here, it is where hydrodynamical simulations (e.g., Schaye et al. 2014; Vogelsberger et al. 2014; Crain et al. 2015; Nelson et al. 2019; Wetzel et al. 2023) can play a crucial role, offering a powerful framework to explore galaxy formation in a cosmological context and validating the methods typically used to infer the internal structure of galaxies, including their dark matter content. One such example is the validation of the Jeans Anisotropic Modelling (JAM for short, e.g., Cappellari et al. 2013a; Cappellari 2016), which is one typical manner to investigate the mass distribution of galaxies by combining two-dimensional kinematic information with dynamical modelling. By applying the Multi-Gaussian Expansion (MGE) formalism (Emsellem et al., 1994) to the projected light distribution of galaxies from the Illustris simulations (Vogelsberger et al., 2014), it was demonstrated by Li et al. (2015) that the JAM approach is a high accuracy method, being able to reproduce the distribution of mass in galaxies with negligible bias.
Unfortunately, techniques involving dynamical modelling, such as JAM, often need rather long computational times. A consequence is that such approaches are more feasible for samples of thousands of galaxies (see, e.g., Tortora et al. 2014; Tortora & Napolitano 2022; Zhu et al. 2023b) while the upcoming large galaxy surveys will observe millions to billions of galaxies. In this context, machine Learning (ML) models present a promising avenue for addressing this challenge. One example of its use is as an alternative approach to inference of the total and dark matter content in galaxies (e.g., Von Marttens et al. 2022; Wu et al. 2024; Chu et al. 2024) or the emulation of dynamical models such as JAM with neural networks (Gomer et al., 2023). In particular, in Wu et al. (2024), we have provided a first application of these techniques to real samples and shown that machine learning can be as accurate as JAM or similar Jeans analyses, but providing total and dark masses of thousands of galaxies in a timescale of a few seconds.
@@ 2.1 The simulation
The galaxies are identified as gravitationally bound structures of stellar, gas and dark matter particles using the SUBFIND algorithm (Springel et al., 2001; Dolag et al., 2009). For this study, we selected only central galaxies that are the most massive objects identified by SUBFIND in each friend-of-friend (FoF) group and are of cosmological origin (SubhaloFlag=True), in the sense that they have formed due to the process of structure formation and collapse (Nelson et al., 2019). We filtered galaxies in nine snapshots corresponding to the redshift values $z=0,0.01,0.02,0.03,0.05,0.06,0.07,0.08,0.1$. We assume that in this redshift window there is no significant galactic evolution to impact the studies of this paper and, for now, we do not try to inspect the impact of redshift in the analysis. We also consider only the galaxies with total stellar masses (extracted from the SUBFIND catalogue) in the range $10^{10.3}\leq M_{\star}/M_{\odot}\leq 10^{11.9}$, to guarantee a sufficiently large number of stellar particles per galaxy to minimize the Poissonian noise on the average quantities derived by the binning of the particles (see below). This first selection results in a sample of $23161$ objects for $z\leq 0.1$. Further filters to this sample (such as galaxy morphological classification) are discussed in detail in the next sections.
@@ 2.3.1 Stellar light measurement
where $\tau$ is the optical depth. In this model, the $\tau$ for young stellar systems is associated with dust in molecular clouds, in which the young stars are embedded. Following the disruption of molecular clouds, which occurs on a timescale of $t\sim 10^{7}$ yr (Blitz & Shu, 1980), the optical depth is associated with a uniform screen across the galaxy. Following Charlot & Fall (2000), we use $\tau_{1}\sim 1.0$ and $\tau_{2}\sim 0.3$, that have been found to adequately describe a wide range of observations.
@@ 2.3.2 Galaxy geometry: ellipticity, centroid of luminosity, and orientation angle
Before we move to detail this procedure, it is important to recall that when a smaller halo is accreted onto a bigger structure, its dark and baryonic matter at the outskirts can be tidally stripped while sinking to the centre of the host. This stellar mass is still “weakly” gravitationally bound to the galaxy, and hence normally included in the raw measurement of SUBFIND. In the context of galaxy clusters, this stellar material is usually referred as intracluster light, and it generally accounts for a significant fraction of the total luminosity of a galaxy it overlaps with (Lin & Mohr, 2004; Zibetti et al., 2005; Puchwein et al., 2010). Despite its importance in the context of the galaxy and group/cluster formation theories, this component represents a contaminant of the individual galaxies, as it does not constitute a fully bound component of any specific individual galaxy, but rather a component eventually relaxing in the cluster potential (see, e.g., Pota et al. 2018; Spiniello et al. 2018), with a phase space distribution of the stars which is different from the individual galaxy populations. To mitigate the impact of this component on the stellar mass and derived light profiles, we applied a two-step selection process, following Xu et al. (2017). First, we employed a commonly used technique: a 3D radial cut at 30 kpc from the galactic centre (e.g., Pillepich et al. 2017; Lu et al. 2020). The galactic centre is defined as the position of the particle with the minimum gravitational potential, as determined by the SUBFIND algorithm. Next, the selected stellar particles were projected onto a 2D plane, where we defined the galaxy’s centroid of light $(X_{gc},Y_{gc})$. This centre is computed as the luminosity-weighted average position of the projected stellar particles:
@@
where $L_{i}$ represents the dust-attenuated luminosity of each stellar particle. Finally, disregarding the previous 3D cut, we applied a 2D radial cut that selects all projected stellar particles within a radius of 30 kpc from $(X_{gc},Y_{gc})$.
Given the values of $(X_{gc},Y_{gc})$, the measurement of the axis ratio $K=b/a$ and orientation angle $\phi$ of a galaxy were finally obtained through the computation of the second moments of luminosity, given by the set of equations:
@@ 2.3.3 Surface brightness profiles
As for the maximum radius to adopt to include stellar particles in the fit, we have used a $30$ kpc aperture (see, e.g., Xu et al. 2017; Pillepich et al. 2018) centred at $(X_{gc},Y_{gc})$, according to the previous discussion. We have also tried a more observationally motivated SB cut, corresponding to 26.5 mag/arcsec2 in $g$-band (see Tang et al. 2021, and also §1). However, since we have found no significant differences in the best-fit quantities that we will discuss later in this paper (in particular the effective radius and the total luminosity of the virtual-ETG sample), we have decided to discard the quantities based on the SB cut. The reason for the low sensitivity to the outer cut resides on the fact that this does not directly impact the inferred quantities that are always the result of a fitting procedure. This is different from the reason motivating the 26.5 mag/arcsec2 in $g$-band in Tang et al. (2021), who use as total quantities the ones obtained by integrating all particles within that limit, with neither fit or extrapolation.
@@
We chose the best-fit parameters based on a minimum $\chi^{2}$ approach, having used $R_{\rm e,dir}$ and $I_{\rm e,dir}$ for each galaxy as initial guesses for the fitting procedure, while $n=2.5$ is assumed as initial values for all galaxies. Using a least-squares algorithm [^6] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.least_squares.html, we fitted several profiles considering the different allowed values of $x$ (with $5$ steps between each sub-range) and chose the profile in which the fitted parameters give the minimum value of $\chi^{2}$. In short, ten profiles with different lower radial limits were fitted to each galaxy, and the best-fit was chosen based on the minimum $\chi^{2}$ value among them. After the fitting step, we took $R_{\rm e,mod}$, $I_{\rm e,mod}$ and $n$ as being the final half-light radius, half-light surface brightness and Sérsic index for each galaxy. For these ones, from now on, we drop the subscript mod, and refer to them as $R_{\rm e}$, $I_{\rm e}$ and $n$, respectively. Particular values of $n$ in Eq. (9) include $n=1$ that is the so-called exponential profile (Mo et al., 2010; Binney & Tremaine, 2011), and characterizes the surface brightness distribution of disky or other light-diffuse systems. For $n=4$ we have the De Vaucouleurs profile (De Vaucouleurs, 1948; Mo et al., 2010; Binney & Tremaine, 2011), that is better suited for the description of spheroidal systems, such as bulges and elliptical galaxies. In Fig. 1, we show the visual representations of two example galaxies and their respective projections. The one in the upper row has a low Sérsic index value ($n\lesssim 1$), and a disky appearance, which characterizes a Late-Type Galaxy (LTG). The galaxy in the lower row has a higher Sérsic index ($n\geq 4$) and has a smooth, featureless appearance, which are characteristics of an ETG. Furthermore, in Fig. 2 we present the fitted surface brightness profiles and their mock measruments for a pair of ETGs and a pair of LTGs, projected along the simulation’s $X$-axis [^7] While we initially omitted uncertainties in the light profiles, we acknowledge that they could be estimated from the brightness scatter within each bin. However, our tests show that including these does not significantly affect the Sérsic fits or the scaling relations examined (see §3), and thus does not impact our discussion.. More details on the galaxy morphological classification are presented in §2.3.4.
@@ 2.6.1 Three-dimensional quantities
    Total 3D stellar mass, ($M_{\rm 3D,\star}^{\rm tot}$): this is simply derived as $M_{\rm 3D,\star}^{\rm tot}=M_{\rm tot}-M_{\rm DM}-M_{\rm gas}$. All these quantities are derived from the TNG100 SUBFIND catalogue888 For a projected total stellar mass definition, a straightforward approach is given by Eq. (12), that gives a logical way of estimating the modelled luminosity of a galaxy at any radius (Trujillo et al., 2001). Formally, the total luminosity of a galaxy is defined at an infinite extent, $r\to\infty$ in Eq. (12). Based on this total luminosity definition, the total stellar mass of a galaxy can be approximated as $M_{\star}(r\to\infty)=\Upsilon_{\star}L(r\to\infty)$. One should not expect this stellar mass definition to equalize the total stellar mass $M_{\rm 3D,\star}(<30~{}\text{kpc})$, as the latter is truncated at a finite radius while the former allows for an analytical extrapolation up to infinity. In fact, based on this extrapolation, $M_{\rm 2D,\star}^{\rm tot}(<\infty)$ should be greater than $M_{\rm 3D,\star}(<30~{}\text{kpc})$, specially at the bright end, where we would expect a greater impact of the intracluster light. With that in mind, we only discuss the comparison between the aperture luminosity and stellar mass definitions and their projected counterparts in §3.4.2 and §3.4.3..
@@ 2.6.2 Projected quantities
    where $\bar{v}$ is the mean velocity of the same $v_{i}$ sample defined above  [^10] We have also computed a light-weighted velocity dispersion, using the ratio of the particle luminosity over the total luminosity as weight, obtaining the same distribution of $\sigma_{[\epsilon_{\star},R_{e}]}$, consistently with the recent findings from Sohn et al. (2024), hence we decided to keep the simple Eq. (19).. This is the most unbiased LOS velocity dispersion we can “measure” in simulations that fully traces the internal dynamics of the simulated galaxies and that is closely connected to the “virial” velocity dispersion that one should include in the galaxy fundamental plane (de Graaff et al. 2022), although in Eq. (19) we include both random and rotational motions making this definition intrinsically larger than the rotation subtracted velocity dispersions derived from galaxy spectroscopy, especially toward the low mass end ($\log M_{\star}/M_{\odot}<11$) of our sample that is usually dominated by fast-rotators (see, e.g., Guo et al. 2020). Furthermore, real spectroscopy, yet differently from the $\sigma_{[\epsilon_{\star},R_{\rm e}]}$ above, considers all stars down to the very galaxy centres, where the velocity dispersion profile of ETGs, usually show their maximum value111111 For observations, the main “dilution” factor is the atmospheric seeing, which is generally ignored on fiber/long-slit observations, but it is accounted for in integral field spectroscopy (Chung et al. 2021 and references therein).. To take this central component, missed by Eq. (19), into account, we introduce a “corrected” $\sigma_{\rm e}$, which is fully detailed in Appendix D but that we also report here below for sake of completeness:
@@
    This $\sigma_{\rm e,corr}$ is the one we will use in the analysis and comparisons with literature, and will be included in the catalogue. Thus, from now on we will drop the subscript $\rm corr$ and refer to it just as $\sigma_{\rm e}$, unless it is needed to be specified.